Matt Johnson delivers newspapers and is putting away ​$17 at the end of each month from his paper route collections. Matt is 9 years old and will use the money when he goes to college in 9 years. What will be the value of​ Matt's account in 9 years with his monthly payments if he is earning 5​% (APR), 9%(APR), or 13.5% (APR)?

What will be the value of​ Matt's account in 9 years with his monthly payments if he is earning 5​% (APR)?

​(Round to the nearest​ cent.)
Answer #1

Calculating Future Value,

At APR = 5%

FV = [PV = 0, PMT = 17, N = 108, I = 0.05/12]

FV = $2,312.73

At APR = 9%

FV = [PV = 0, PMT = 17, N = 108, I = 0.09/12]

FV = $2,813.21

At APR = 13.5%

FV = [PV = 0, PMT = 17, N = 108, I = 0.13.5/12]

FV = $3,547.35